OpenStudy (anonymous):

What is the slope of the line that contains the points (-8, 7) and (-4, -5)? -3 1/3 3 -1/3

OpenStudy (anonymous):

@texaschic101 could you help me with thisone it ismylast

OpenStudy (anonymous):

@nuccioreggie @imqwerty @steve816 @skullpatrol

OpenStudy (texaschic101):

ok...we are going to use the slope formula again.. slope = (y2 - y1) / (x2 - x1) we have 2 sets of points... set 1 : (-8,7)....so we will call these : x1 = -8 and y1 = 7 set 2 : (-4,-5)...and we will call these : x2 = -4 and y2 = -5 now we sub slope = (-5-7) / (-4-(-8) slope = (-12) / (-4 + 8) slope = -12/4 slope = - 3 any questions ?

OpenStudy (anonymous):

ohhhhhhh i getit now,.. sorry lol andt hank you and no im gutchi

OpenStudy (texaschic101):

good to hear :) always glad to help out